HETAV PATEL

Grade 11,

in hydrogen atom an electron of mass 9.1x10^-31kg revolves about a proton in a circular orbit of radius 0.53A0.the centrifugal acceleration and angular velocity of the electron is as??

in hydrogen atom an electron of mass 9.1x10^-31kg revolves about a proton in a circular orbit of radius 0.53A0.the centrifugal acceleration and angular velocity of the electron is as??
 

Grade:12

1 Answers

Vikas TU
7 years ago
Centriputal acceleration that is,
From Bohr’s angular momentum concept,
mvr = h/2pi
9.1x10^-31*v*0.53*10^-10 = h/2pi
v = 1.057*10^-34/9.1*10^-31*0.53) = 2.19 x 10^-4 m/s
a = v^2/r => (4*10^-8)/(0.53*10^-10) = 887.76 m/s^2
w = v/r => 2.19 x 10^-4/(0.53*10^-10) = 4132075 rad/s